# Graph-Based Approximate Nearest Neighbor Search Revisited: Theoretical Analysis and Optimization

### v2

@@ 6.2 Search Performance
Figure 3: Query performance across varying $\alpha$ values on GIST1M
Figure 4: Query performance across varying $\alpha$ values on DEEP1M
6.3 Scalability Evaluation
To evaluate the scalability of our theoretical results—particularly the sparsity and degree bounds established under uniform distributions in Section 3—we extend our analysis to non-uniform data settings. Specifically, we generate a synthetic dataset consisting of n=100,000n=100{,}000 points drawn from a Gaussian Mixture Model (GMM) with 8 dimensions and 10 clusters. We use 90% of the data (90,000 points) as the base set to construct an SNG under parameter $\alpha=1.2$. Results under other values of $\alpha$ are also provided in Appendix H.4.
Figure 5: Degree distribution of GMM dataset
The degree distribution of all nodes is visualized in Figure 5. Notably, approximately 20,000 points have degrees concentrated in the range $[50,60]$, and the maximum observed degree does not exceed 100. This empirical evidence supports our theoretical prediction that the graph degree remains bounded by $O(n^{2/3})$, even under complex, non-uniform distributions. The corresponding query performance on this dataset is reported in Table 3.
7 Conclusions
In this paper, we revisited the design and analysis of graph-based ANNS methods, with a focus on SNG). Our work bridges the gap between theoretical understanding and practical deployment. We provide rigorous theoretical guarantees on the structure of SNG by proving that the graph degree is bounded by $O(n^{2/3})$ and that the search path length is $O(\log n)$. Building on this theoretical foundation, we introduced a novel parameter optimization strategy based on the marginal optimality principle. Unlike existing heuristic methods such as binary search, our approach yields a principled estimate for the truncation parameter $R$, significantly reducing the tuning overhead while maintaining or even improving search performance. Experimental results on a GMM dataset further validate the scalability of our theoretical analysis of degree.
@@ Appendix A Supplementary Lemmas
The following Chernorff bound provides a probabilistic guarantee on the deviation of a binomial random variable from its expected value.
@@
The frame of Wormald’s differential equation method (DEM) was developed by Wormald in 1990s [35] as a powerful tool for analyzing the discrete-time randomized graph processes and algorithms. Given a discrete-time stochastic process, in particular, consider a submartingale, if the one-step differences are bounded and the expected differences are well-approximated by a Lipschitz function, it follows that the process closely aligns with the corresponding differential equation with high probability.
Lemma 11 (Differential Equation method [35]).
@@
Appendix B Summary of Notations
The following is a summary of the main symbols used in the paper.
Table 1: Summary of Notation
Symbol Description
$P=\{p_{1},\ldots,p_{n}\}$ Dataset of $n$ points in $\mathbb{R}^{d}$
$S_{t}$ Candidate set at iteration $t$
$p^{*}_{t}$ Nearest neighbor of $p$ in $S_{t}$ at iteration $t$
$\mathcal{F}_{t}$ Natural filtration up to iteration $t$
$\Delta S_{t}$ Number of points pruned in iteration $t$
$\pi_{t}$ Pruning probability at iteration $t$
$R$ Truncation parameter (maximum degree limit)
$M$ Number of processed points at a given level
$t$ Iteration index (time step)
Appendix C Omitted Proofs
C.1 Proof of Lemma 6
@@
Lemma 6.
Given a dataset $P\subset\mathbb{R}^{d}$ consisting of $n$ points independently and uniformly distributed within a ball of radius $\rho_{0}$, centered at $p$. we consider the SNG construction processed on the center point $p$ with $\alpha=1$. The probability $\pi_{t}$ that a point $p^{\prime}$ is pruned in the $t$-th iteration is given by $\pi_{t}=\frac{I_{1-(\rho_{t}/\rho_{0})^{2}}\left(\frac{d+1}{2},\frac{1}{2}\right)-(\rho_{t}/\rho_{0})^{d}\cdot I_{0.75}\left(\frac{d+1}{2},\frac{1}{2}\right)}{2\left(1-(\rho_{t}/\rho_{0})^{d}\right)},$ where $I_{x}(a,b)$ denotes the regularized incomplete Beta function, and $\rho_{t}$ is the distance between the nearest neighbor and $p$ in ttht-th iteration. For large $d$, this probability satisfies $\frac{I_{0.75}\left(\frac{d+1}{2},\frac{1}{2}\right)}{2}<\pi_{t}<\frac{1}{2}.$
@@
Figure 6: Given that $p^{*}$ is the nearest point to $p$ in $t$-th iteration and $\|p-p^{*}\|=\rho_{t}$. The light grey region represents all $p^{\prime}$ satisfied pruning condition $\|p-p^{\prime}\|>\|p^{*}-p^{\prime}\|$.
@@
C.2 Conservative Probability Estimation
@@
C.3 Proof of Lemma 8
Proof of Lemma 8.
Part (i): We aim to show that $|S^{\prime}_{t_{1}}|\geq n^{\nu_{1}}$ for $t_{1}=O(n^{\nu_{2}})$. To that end, define the first passage time $t_{1}$ as the smallest $t$ such that $|S^{\prime}_{t}|\geq n^{\nu_{1}}$. Our goal to prove that this event occurs with probability one, i.e.,
$\Pr(t_{1}\leq Cn^{\nu_{2}})=1$
for some $C>0$. Formally:
$\Pr(t_{1}=O(n^{\nu_{2}}))=\Pr\left(\bigcup_{C>0}\bigcup_{n_{0}>0}\bigcap_{n>n_{0}}\{t_{1}\leq Cn^{\nu_{2}}\}\right)=\Pr\left(\bigcup_{C>0}\bigcup_{n_{0}>0}\bigcap_{n>n_{0}}\{|S^{\prime}_{Cn^{\nu_{2}}}|\geq n^{\nu_{1}}\}\right).$
This is equivalent to demonstrating that:
$\Pr\left(\bigcap_{C>0}\bigcap_{n_{0}>0}\bigcup_{n>n_{0}}\{|S^{\prime}_{Cn^{\nu_{2}}}|<n^{\nu_{1}}\}\right)=0.$
Now, consider the probability $\Pr(|S^{\prime}_{Cn^{\nu_{2}}}|<n^{\nu_{1}})$. Since $|S^{\prime}_{t}|=\sum_{i=1}^{t}(\Delta S_{i}+1)$, for $t=Cn^{\nu_{2}}$, we have:
$|S^{\prime}_{Cn^{\nu_{2}}}|=\sum_{i=1}^{Cn^{\nu_{2}}}(\Delta S_{i}+1)\geq Cn^{\nu_{2}}\cdot\min_{1\leq i\leq Cn^{\nu_{2}}}(\Delta S_{i}+1).$
Thus, we have:
$\Pr(|S^{\prime}_{Cn^{\nu_{2}}}|<n^{\nu_{1}})$ $\leq\Pr\left(Cn^{\nu_{2}}\cdot\min_{1\leq i\leq Cn^{\nu_{2}}}(\Delta S_{i}+1)<n^{\nu_{1}}\right)$
$=\Pr\left(\min_{1\leq i\leq Cn^{\nu_{2}}}\Delta S_{i}<\frac{n^{\nu_{1}}-Cn^{\nu_{2}}}{Cn^{\nu_{2}}}\right).$
Since $\nu_{1}>\nu_{2}$, for large $n$, we have $n^{\nu_{1}}\gg Cn^{\nu_{2}}$, which implies that $\frac{n^{\nu_{1}}-Cn^{\nu_{2}}}{Cn^{\nu_{2}}}\approx\frac{n^{\nu_{1}}}{Cn^{\nu_{2}}}$. We apply the Chernoff bound (Lemma 10) to estimate $\Pr(\Delta S_{i}\leq(1-\delta)E[\Delta S_{i}])$:
$\Pr(\Delta S_{i}\leq(1-\delta)E[\Delta S_{i}])\leq\exp\left(-\frac{\delta^{2}E[\Delta S_{i}]}{2}\right)\leq\exp\left(-\frac{\delta^{2}(|S_{i-1}|-1)I_{0.75}\left(\frac{d+1}{2},\frac{1}{2}\right)}{4}\right),$
where $E[\Delta S_{i}]=\pi_{i}(|S_{i-1}|-1)$ and $\pi_{i}>\frac{I_{0.75}\left(\frac{d+1}{2},\frac{1}{2}\right)}{2}$. We set $1-\delta=\frac{\frac{n^{\nu_{1}}-Cn^{\nu_{2}}}{Cn^{\nu_{2}}}}{E[\Delta S_{i}]}\geq\frac{\frac{n^{\nu_{1}}-Cn^{\nu_{2}}}{Cn^{\nu_{2}}}}{n/2}$, noting that $E[\Delta S_{i}]\leq|S_{i-1}|-1\approx n$ early in the process. Thus, we obtain:
$\Pr(|S^{\prime}_{Cn^{\nu_{2}}}|<n^{\nu_{1}})\leq\exp\left(-\frac{\left(1-\frac{\frac{n^{\nu_{1}}-Cn^{\nu_{2}}}{Cn^{\nu_{2}}}}{n/2}\right)^{2}(n-n^{\nu_{2}}-1)I_{0.75}\left(\frac{d+1}{2},\frac{1}{2}\right)}{4}\right).$
The series $\sum_{n>0}\exp(-\frac{\left(1-\frac{\frac{n^{\nu_{1}}-Cn^{\nu_{2}}}{Cn^{\nu_{2}}}}{n/2}\right)^{2}(n-n^{\nu_{2}}-1)I_{0.75}\left(\frac{d+1}{2},\frac{1}{2}\right)}{4})$ converges, as the exponent becomes increasingly negative with $n$. By the Borel-Cantelli lemma (Lemma 9), we have $Pr\left(\bigcap_{C>0}\bigcap_{n_{0}>0}\bigcup_{n>n_{0}}\{|S^{\prime}_{Cn^{\nu_{2}}}|<n^{\nu_{1}}\}\right)=0$, which implies that $\Pr(t_{1}=O(n^{\nu_{2}}))=1$.
Part (ii): The proof follows a similar approach. We need to demonstrate that $|S^{\prime}_{t_{2}}|\geq n-n^{1-\nu}$ for $t_{2}=O(n^{\nu})$. Let $t_{2}$ be the first passage such that $|S^{\prime}_{t}|\geq n-n^{1-\nu}$. We compute:
$\Pr(t_{2}=O(n^{\nu}))=\Pr\left(\bigcup_{C>0}\bigcup_{n_{0}>0}\bigcap_{n>n_{0}}\{|S^{\prime}_{Cn^{\nu}}|\geq n-n^{1-\nu}\}\right).$
Similarly, we have:
$\Pr(|S^{\prime}_{Cn^{\nu}}|<n-n^{1-\nu})\leq\Pr\left(\min_{1\leq i\leq Cn^{\nu}}\Delta S_{i}<\frac{n-n^{1-\nu}-Cn^{\nu}}{Cn^{\nu}}\right).$
Let $1-\delta\geq\frac{\frac{n-n^{1-\nu}-Cn^{\nu}}{Cn^{\nu}}}{n/2}$. The Chernoff bound provides a convergent series:
$\sum_{n>0}\exp\left(-\frac{\left(1-\frac{\frac{n-n^{1-\nu}-Cn^{\nu}}{Cn^{\nu}}}{n/2}\right)^{2}(n-n^{\nu}-1)I_{0.75}\left(\frac{d+1}{2},\frac{1}{2}\right)}{4}\right)<\infty.$
By the Borel-Cantelli lemma, $\Pr\left(\bigcap_{C>0}\bigcap_{n_{0}>0}\bigcup_{n>n_{0}}\{|S^{\prime}_{Cn^{\nu}}|<n-n^{1-\nu}\}\right)=0$, which implies $\Pr(t_{2}=O(n^{\nu}))=1$. ∎
Part (i) indicates that the initial phase of SNG construction exhibits rapid growth, with the number of processed points scaling exponentially faster than the number of iterations. In contrast, part (ii) demonstrates that a sublinear number of iterations, on the order of $O(n^{\nu})$, is sufficient to process nearly all points, achieving a growth magnitude close to $n$.
C.4 Proof of Theorem 1
@@
First Phase ($t\leq t_{1}$): From Lemma 8 (part ii), for any $\nu\in(0,1)$, with probability 1, the SNG construction reaches the $(n-n^{1-\nu})$-$O(n^{\nu})$ level. Thus, the number of iterations $t_{1}$ to process $n-n^{1-\nu}$ points is:
@@
$|S^{\prime}_{t+1}|-|S^{\prime}_{t}|=\Delta S_{t+1}+1,\quad E[|S^{\prime}_{t+1}|-|S^{\prime}_{t}|\mid\mathcal{G}_{t}]=1+\pi_{t}(|S_{t}|-1),$
where $\mathcal{F}_{t}$ is the natural filtration up to iteration $t$ (defined in Section 2.2). Let $z(t/n):=|S^{\prime}_{t}|/n$ be the normalized process, and approximate the expected increment by:
@@
This verifies the Lipschitz condition for the application of Lemma 11. The corresponding difefrential equation is:
@@
where $\rho_{0}$ is the radius of the ball, $h$ is the distance between centers, and $I_{x}(a,b)$ denotes the regularized incomplete Beta function, with the detailed derivation provided in [23]. Due to the computability of this volume, the pruning probability can be updated and remains bounded. The proof technique can be adapted accordingly, thus preserving the conclusions regarding the degree bound. ∎
C.5 Full Proof to Theorem 2
@@
Figure 7: Path and neighbor distribution diagram: Blue lines represent edges from $v_{k}$ to its neighbors, and green lines represent the search path. $\pi_{i}$ denotes the number of neighbors in the $i$-th annular layer from the outermost to the innermost.
@@
$\triangle r=\min\{|\|a-b\|-\|a-c\||,|\|a-b\|-\|b-c\||,|\|a-c\|-\|b-c\||\mid a,b,c\in P,\text{distinct}\}$
as introduced in [12]. Let $v_{k}$ be the final node of the search path. We construct concentric balls centered at $v_{k}$ with radii $\|v_{k}-v_{i}\|$ for i=0,1,,k1i=0,1,...,k-1, and define the annular layer $A_{i}=B(v_{k},\|v_{k}-v_{i-1}\|)\setminus B(v_{k},\|v_{k}-v_{i}\|)$. Let $\eta_{i}$ be the number of neighbors in each layer $A_{i}$. Then $\eta_{i}\sim\text{Bin}\left(\eta,\frac{V(A_{i})}{V(B(v_{k},\rho_{0}))}\right)$, and the expected number in each layer is proportional to the volume difference:
@@
$E\left[\sum_{i=1}^{k}\eta\cdot\frac{V(B(v_{k},\|v_{k}-v_{i-1}\|))-V(B(v_{k},\|v_{k}-v_{i}\|))}{V(B(v_{k},R))}\right]\leq n\cdot\frac{V(B(v_{k},\|p-v_{k}\|))}{V(B(v_{k},R))}.$
@@
$E\left[\sum_{i=1}^{k}\eta\cdot\frac{\|v_{k}-v_{i-1}\|^{d}-\|v_{k}-v_{i}\|^{d}}{\rho_{0}^{d}}\right]\leq n\cdot E\left[\frac{\|p-v_{k}\|^{d}}{\rho_{0}^{d}}\right].$ (1)
@@
$E[k\cdot\eta]\leq n\cdot\frac{E[\|p-v_{k}\|^{d}]}{\triangle r\cdot d\cdot E[\|v_{k}-v_{k-1}\|^{d-1}]}.$ (2)
@@
Using the bound $\mathbb{E}[\|v_{k}-v_{k-1}\|]\geq\frac{\rho_{0}}{(n+1)^{1/d}}$ from [29] and Jensen’s inequality(Lemma 12), we find:
@@
Appendix D Foundations of Probability Spaces
In this section, we provide formal definitions of the foundational concepts used in our probabilistic analysis, including $\sigma$-algebras, probability spaces, random variables, and adapted processes. For more detailed content related to probability, refer to [2].
Definition 13 ($\sigma$-algebra).
$\Omega$ is the sample space. A nonempty class $\mathcal{F}$ of subsets of $\Omega$ is a $\sigma$-algebra on $\Omega$ if
(i). $A^{c}\in\mathcal{F}$ whenever $A\in\mathcal{F}$,
(ii). $\cup_{n=1}^{\infty}A_{n}\in\mathcal{F}$ whenever $A_{n}\in\mathcal{F},n\geq 1$.
(i.e. $\mathcal{F}$ is closed under complement and countable union.) The pair ( $\Omega,\mathcal{F}$ ) is called a measurable space. The sets of $\mathcal{F}$ are called measurable sets.
Definition 14 (Probability Space).
A probability space is a triple $(\Omega,\mathcal{F},\mathbb{P})$, where:
(i). $\Omega$ is the sample space (set of all possible outcomes).
(ii). $\mathcal{F}$ is a $\sigma$-algebra over $\Omega$.
(iii). $\mathbb{P}:\mathcal{F}\to[0,1]$ is a probability measure.
Definition 15 (Random Variable).
$(\Omega,\mathcal{F},\mathbb{P})$ is a probability space. A random variable $X$ is a measurable function $X:\Omega\to\mathbb{R}$, meaning that for any Borel set $B\subseteq\mathbb{R}$, the preimage $X^{-1}(B)\in\mathcal{F}$. This ensures that we can assign probabilities to events of the form $X\in B$.
Definition 16 (Adapted Process).
Let $\{\mathcal{F}_{t}\}_{t\geq 0}$ be a filtration (i.e., an increasing sequence of $\sigma$-algebras) on $(\Omega,\mathcal{F},\mathbb{P})$. A stochastic process $\{X_{t}\}_{t\geq 0}$ is said to be adapted to the filtration $\{\mathcal{F}_{t}\}$ if, for every $t$, the random variable $X_{t}$ is $\mathcal{F}_{t}$-measurable. This means that the value of $X_{t}$ is known given the information available up to time $t$.
Appendix E Algorithms
Below is the detailed RobustPrune algorithm used to do truncated SNG pruning for points in the datasets  [30], corresponding to the graph construction process in Section 2.1.
Algorithm 2 RobustPrune($p$, $S$, $\alpha$, $R$)
0: Graph $G$, point $p\in P$, candidate set $S$, relaxation parameter $\alpha$, truncation parameter $R$
0:$G$ with neighbor set $N_{\text{out}}(p)$
1: Set $N_{\text{out}}(p)\leftarrow\emptyset$
2:while $S\neq\emptyset$ do
3:  $p^{*}\leftarrow\arg\min_{s\in S}\|s-p\|)$
4:  Add $p^{*}$ to $N_{\text{out}}(p)$
5:  Remove $p^{*}$ from $S$
6:  if $|N_{\text{out}}(p)|=R$ then
7:   break
8:  end if
9:  for $p^{\prime}\in S$ do
10:   if $\|p-p^{\prime}\|\geq\alpha\cdot\|p^{*}-p^{\prime}\|$ then
11:    Remove $p^{\prime}$ from $S$
12:   end if
13:  end for
14:end while
15:return $G$ with the neighbor points of $p$
Below, we present the overall construction algorithm—Vamana—which consists of two pruning steps and one search step. The algorithm iterates through all points $p\in P$ in a random order to construct the graph. Specifically, in line 11 of the Vamana algorithm, reverse edges are added to enhance the graph’s connectivity.
Algorithm 3 Vamana Algorithm
0: Database $P$ with $n$ points where $i$-th point has coordinates $x_{i}$, parameters $\alpha$, truncation parameter $R$
0: Directed graph $G$ over $P$ with out-degree $\leq R$
1: Randomly initialize $G$ to a random $R$-regular directed graph
2: Let $s$ denote the medoid of dataset $P$
3: Let $\sigma$ denote a random permutation of $1,2,\dots,n$
4:for $i=1$ to $n$ do
5:  Perform a GreedySearch from $s$ to $x_{\sigma(i)}$, $\mathcal{V}$ $\leftarrow$ set of all points visited along the search path
6:  Run $\text{RobustPrune}(\sigma(i),\mathcal{V},\alpha,R)$ using $\mathcal{V}$ as the candidate set
7:  for all points $j\in N_{\text{out}}(\sigma(i))$ do
8:   if $|N_{\text{out}}(j)\cup\{\sigma(i)\}|>R$ then
9:    Run $\text{RobustPrune}(j,N_{\text{out}}(j)\cup\{\sigma(i)\},\alpha,R)$
10:   else
11:    Update $N_{\text{out}}(j)\leftarrow N_{\text{out}}(j)\cup\{\sigma(i)\}$
12:   end if
13:  end for
14:end for
Appendix F Complexity Analysis Details

### v3

@@ 7.4. Discussion on UNIFORM: A Worst-Case Construction Regime
Figure 4. Query performance across varying $\alpha$ values on UNIFORM
7.5. Effect of Pruning Parameter $\alpha$
Figure 5. Index construction time comparison
Figure 6. Average out-degree comparison across datasets
Figure 7. Scalability of index construction speedup on SIFT10M subsets
Figure 8. Query performance on SIFT10M
Figure 3 illustrates the recall versus QPS curves under different $\alpha$ settings. Across almost all datasets, increasing the pruning parameter $\alpha$ exhibits a consistent trend: for a fixed recall level, the achievable QPS initially increases as $\alpha$ grows, and then either saturates or slightly degrades when $\alpha$ becomes too large. We observe that $\alpha=1.50$ provides the best overall balance across datasets, achieving the highest or near-highest QPS at comparable recall levels in most cases.
The sensitivity to $\alpha$ varies across datasets. High-dimensional datasets such as GIST1M and DEEP1M exhibit more pronounced performance degradation at large $\alpha$ values, as overly sparse graphs struggle to preserve navigability in high-dimensional spaces. In contrast, lower- or medium-dimensional datasets such as SIFT1M, GloVe, and MSong remain robust across a wider range of $\alpha$ values, showing only mild performance variation.
7.6. Scalability Study
Finally, we evaluate the scalability of our approach by varying the dataset size and reporting the index construction time. In addition, we conduct experiments on the SIFT10M dataset by sampling subsets of size 2M, 4M, 6M, and 8M points. As shown in Figure 8, the construction speedups achieved by the OPT variants are consistently preserved across all sampled subsets, and the benefits become increasingly pronounced as the dataset size grows.
Furthermore, as illustrated in Figure 8, the recall–QPS curves produced by OPT consistently match or outperform those obtained via parameter sweeping across all three algorithms, without compromising search quality. These results demonstrate that our optimization strategy scales favorably to large datasets and remains effective in both construction efficiency and query performance at scale.
7.7. Summary
The experimental results provide evidence supporting the effectiveness of the proposed analytical parameter optimization strategy. Across all evaluated datasets and SNG-based algorithms, our method substantially reduces end-to-end index construction time by eliminating expensive iterative parameter sweeping. In particular, tuning overhead is reduced by more than an order of magnitude, leading to overall construction speedups of up to $15.4\times$ and an average speedup of $5.9\times$ across datasets.
Crucially, this acceleration does not come at the cost of query quality. The analytically determined parameters consistently preserve—and in many cases improve the recall and QPS compared to baseline parameter sweeping. Recall@10 is matched or improved across all datasets and algorithms at representative latency targets. The results further demonstrate that the proposed method remains robust under the UNIFORM dataset, which represents a worst-case construction regime aligned with our theoretical assumptions.
Scalability experiments on subsets of SIFT10M confirm that the benefits of analytical tuning persist—and become more pronounced—as dataset size increases, while query-time performance remains competitive or superior to parameter sweeping. Taken together, these results validate that our theory-driven approach provides a principled, scalable, and practical alternative to heuristic tuning pipelines for SNG-based ANNS systems.
8. Related Work
8.1. Approximate Nearest Neighbor Search
Over the years, ANNS has evolved to address diverse scenarios, such as multi-metric spaces (Multi), hybrid queries (ACORN), privacy-preserving federated settings (FEDNNsig2024), and retrieval-augmented generation (RAG), which enhances large language models by efficiently querying external knowledge. Methodologically, ANNS techniques are broadly categorized into tree-based (tree1; tree2), hashing-based (LSHarshing; hash2; hash3), quantization-based (quantti1), and graph-based methods (Graph1; Graphprob; Graph3), with graph-based approaches standing out due to their superior real-world performance in high-dimensional and large-scale datasets (Graph2024VLDB; Graph1). Additionally, hybrid systems that combine graph-based and quantization-based approaches, such as DiskANN (NEURIPS2019) and SPANN (Graph2024VLDB), have been introduced to handle billion-scale datasets efficiently. Tree-based methods, such as $k$-d trees (KD-TREE), partition the space hierarchically but suffer from the curse of dimensionality in high-dimensional settings. Hashing-based approaches, exemplified by locality-sensitive hashing (LSH) (LSHarshing), map similar points to the same bucket with high probability but often require substantial memory overhead. Quantization-based techniques, including product quantization (PQ) (DBLP:journals/pami/JegouDS11) and optimized product quantization (OPQ) (OPQ), compress vectors to enable efficient distance computation at the cost of some precision loss.
Among these paradigms, graph-based approaches have emerged as the dominant methodology due to their superior performance on high-dimensional, large-scale datasets (Graph2024VLDB; Graph1; Wang2021ACS). Wang et al. (Wang2021ACS) provide a comprehensive survey and experimental comparison of 13 representative graph-based ANNS algorithms, establishing a unified taxonomy based on seed selection, neighbor selection, and search strategies.
8.2. Graph-Based ANNS Framework
Graph-based ANNS methods have gained significant attention in recent years due to their outstanding performance in high-dimensional spaces and their scalability to large datasets. The state-of-the-art graph-based systems, such as HNSW (HNSW2020), NSG (NSG19), KGraph (Kgraph), DiskANN (NEURIPS2019), DPG (DPG), and SPTAG (SPANN2021) construct graphs with sparse connections to optimize search efficiency. Extensions like Filtered-DiskANN (gollapudi2023filtered) and FreshDiskANN (FRESH) further adapt the framework for different large-scale scenarios. In recent years, graph-based ANNS has also been extended to diverse application settings, including cross-modal retrieval (chen2024roargraph), and filtered search (wu2022hqann; ACORN; ait2025rwalks).
Existing theoretical studies have primarily focused on analyzing fundamental structural properties such as degree bounds, query complexity, and search path length, often under simplified distributional assumptions.
For instance, Laarhoven (ICML2020) establishes sublinear query complexity for threshold-based nearest neighbor graphs under spherical data models. In such graphs, two nodes are connected by an edge if and only if the distance between them falls below a specified threshold value, ensuring that each node maintains connections to all sufficiently close neighbors. This structural property enables efficient greedy routing when the graph captures adequate local connectivity.
Along similar lines, Prokhorenkova and Shekhovtsov (Haya24) analyze the degree distribution of navigable graphs constructed under a binomial distribution. A graph is said to be navigable if a greedy search algorithm can successfully find a path from any starting vertex to any target vertex in the dataset without getting trapped in local minima. Their analysis demonstrates that navigable graphs can achieve average degree $O(n^{\beta})$ for $\beta<1/2$ while preserving the navigability property, revealing a fundamental trade-off between graph sparsity and search reliability.
Beyond degree bounds, another important line of work concerns structural properties that guarantee efficient search convergence. The Monotonic Search Network (MSNET) (NSG19) is a navigable graph with the additional constraint that the search path between any two nodes follows a monotonically decreasing distance trajectory toward the target. This monotonicity property ensures that greedy search makes consistent progress at each step, thereby guaranteeing logarithmic search path length.
More recently, (revisiting2025) proposes a new construction framework incorporating a self-iterative strategy for edge selection. This approach accelerates RNG construction while preserving competitive $k$-ANN search performance, demonstrating that careful algorithmic design can significantly reduce preprocessing overhead without sacrificing retrieval quality.
8.3. Parameter Tuning for ANNS Systems
Selecting optimal parameters for ANNS systems remains a practical challenge. Traditional approaches rely on parameter sweeping over the parameter space, which can be computationally expensive due to repeated index construction and evaluation. Recent work on automated tuning, such as VDTuner (vdtuner), addresses this challenge through learning-based approaches that predict optimal configurations. Our work takes a complementary approach by deriving analytical expressions for parameter selection based on theoretical insights into the pruning process, thereby eliminating the need for iterative search.
9. Conclusions
This paper revisits SNG-based graph indexing for approximate nearest neighbor search from a theoretical and optimization perspective. We develop a martingale-based probabilistic framework, OPT-SNG, to model the stochastic pruning dynamics in SNG construction, which enables a principled analysis of the graph structure generated by widely used systems. We establish that the constructed SNG admits a maximum out-degree of $O(n^{2/3+\epsilon})$ (for any $\epsilon>0$) and that GreedySearch converges in an expected $O(\log n)$ number of steps, providing an expected-case, cardinality-based complement to existing analyses.
Building on these guarantees, we further derive a closed-form rule for selecting the truncation parameter $R$ as a function of $\alpha$ and $n$, and propose an efficient calibration procedure that estimates the scaling constant from a lightweight reference construction. Experiments on multiple real-world benchmarks and a synthetic UNIFORM dataset demonstrate that our analytical parameter selection eliminates expensive parameter sweeping, achieving an average $5.9\times$ reduction in construction time across datasets while maintaining or improving search performance.
We hope this work helps bridge the gap between the empirical success of SNG-based ANNS and its theoretical understanding, and encourages future studies on distribution-aware analyses and principled parameter optimization for large-scale vector search systems.
@@ Appendix A Supplementary Lemmas
The following Chernoff bound provides a probabilistic guarantee on the deviation of a binomial random variable from its expected value.
@@
The frame of Wormald’s differential equation method (DEM) was developed by Wormald in 1990s (DEM1) as a powerful tool for analyzing the discrete-time randomized graph processes and algorithms. Given a discrete-time stochastic process, in particular, consider a submartingale, if the one-step differences are bounded and the expected differences are well-approximated by a Lipschitz function, it follows that the process closely aligns with the corresponding differential equation with high probability.
Lemma 0 (Differential Equation method (DEM1)).
@@
Appendix B Omitted Proofs
B.1. Proof of Lemma 1
@@
Figure 9. Given that $p^{*}$ is the nearest point to $p$ in $t$-th iteration and $\|p-p^{*}\|=\rho_{t}$. The light grey region represents all $p^{\prime}$ satisfied pruning condition $\|p-p^{\prime}\|>\|p^{*}-p^{\prime}\|$.
@@
B.2. Conservative Probability Estimation
@@
B.3. Proof of Theorem 4
@@
First Phase ($t\leq t_{1}$): From Lemma 3 (part ii), for any $\nu\in(0,1)$, with probability one, the SNG construction reaches the $(n-n^{1-\nu})$-$O(n^{\nu})$ level. Thus, the number of iterations $t_{1}$ to process $n-n^{1-\nu}$ points is:
@@
$|S^{\prime}_{t+1}|-|S^{\prime}_{t}|=\Delta S_{t+1}+1,\quad E[|S^{\prime}_{t+1}|-|S^{\prime}_{t}|\mid\mathcal{F}_{t}]=1+\pi_{t}(|S_{t}|-1),$
where $\mathcal{F}_{t}$ is the natural filtration up to iteration $t$ (defined in Section 3). Let $z(t/n):=|S^{\prime}_{t}|/n$ be the normalized process, and approximate the expected increment by:
@@
This verifies the Lipschitz condition for the application of Lemma 3. The corresponding differential equation is:
@@
where $\rho_{0}$ is the radius of the ball, $h$ is the distance between centers, and $I_{x}(a,b)$ denotes the regularized incomplete Beta function, with the detailed derivation provided in (volume). Due to the computability of this volume, the pruning probability can be updated and remains bounded. The proof technique can be adapted accordingly, thus preserving the conclusions regarding the degree bound. ∎
B.4. Full Proof to Theorem 1
@@
Figure 10. Path and neighbor distribution diagram: Blue lines represent edges from $v_{k}$ to its neighbors, and green lines represent the search path. $\pi_{i}$ denotes the number of neighbors in the $i$-th annular layer from the outermost to the innermost.
@@
[…]
$\mid a,b,c\in P,\text{distinct}\}$
as introduced in (NSG19). Let $v_{k}$ be the final node of the search path. We construct concentric balls centered at $v_{k}$ with radii $\|v_{k}-v_{i}\|$ for i=0,1,,k1i=0,1,...,k-1, and define the annular layer $A_{i}=B(v_{k},\|v_{k}-v_{i-1}\|)\setminus B(v_{k},\|v_{k}-v_{i}\|)$. Let $\eta_{i}$ be the number of neighbors in each layer $A_{i}$. Then $\eta_{i}\sim\text{Bin}\left(\eta,\frac{V(A_{i})}{V(B(v_{k},\rho_{0}))}\right)$, and the expected number in each layer is proportional to the volume difference:
@@
(22) $E\left[\sum_{i=1}^{k}\eta\cdot\frac{V(B(v_{k},\|v_{k}-v_{i-1}\|))-V(B(v_{k},\|v_{k}-v_{i}\|))}{V(B(v_{k},R))}\right]$
(23) $\leq n\cdot\frac{V(B(v_{k},\|p-v_{k}\|))}{V(B(v_{k},R))}.$
@@
(24) $E\left[\sum_{i=1}^{k}\eta\cdot\frac{\|v_{k}-v_{i-1}\|^{d}-\|v_{k}-v_{i}\|^{d}}{\rho_{0}^{d}}\right]\leq n\cdot E\left[\frac{\|p-v_{k}\|^{d}}{\rho_{0}^{d}}\right].$
@@
(27) $E[k\cdot\eta]\leq n\cdot\frac{E[\|p-v_{k}\|^{d}]}{\triangle r\cdot d\cdot E[\|v_{k}-v_{k-1}\|^{d-1}]}.$
@@
Using the bound $\mathbb{E}[\|v_{k}-v_{k-1}\|]\geq\frac{\rho_{0}}{(n+1)^{1/d}}$ from (Dist2010) and Jensen’s inequality(Lemma 4), we find:
@@
Appendix C Complexity Analysis Details